# LongForm: Effective Instruction Tuning with Reverse Instructions

### v2

@@ 4.1 LongForm Models
We create three LongForm models by finetuning language models on the LongForm-C dataset to assess their effectiveness on instruction-following text generation: LongForm-T5-3B, based on T5-3B Raffel et al. (2020); Lester et al. (2021), an encoder-decoder LM, and two autoregressive LMs: LongForm-OPT-6.7B, based on OPT-6.7B Zhang et al. (2022), and LongForm-LLaMA-7B, based on LLaMA-7B Touvron et al. (2023).
We include an end-of-instruction token, [EOI], between the instruction and the output for autoregressive LMs, OPT and LLaMA variants, to differentiate between instruction and output. We employ two to four A6000 48GB GPUs for finetuning with a batch size of 32 to 64, and train with DeepSpeed, including bfloat16 training. The learning rate is determined by comparing the minimum validation loss during three epochs: 5e-5 for LongForm-T5-3B, 5e-6 for LongForm-OPT-6.7B and LongForm-LLaMA-7B. Apart from the three large LMs, we finetune four smaller OPT LMs (125M, 350M, 1.3B, 2.7B) with LongForm-C and release them to increase accessibility of instruction-tuned LMs. See §C for more details.
@@ 4.2 Baselines
We compare our instruction-tuned LMs with other LMs that are instruction-tuned on different datasets. These LMs are based on either T5-11B Raffel et al. (2020); Lester et al. (2021) or LLaMA-7B Touvron et al. (2023). Additionally, we compare our LMs with a raw large LMs, OPT-30B Zhang et al. (2022). Notably, all LMs have the same or higher number of parameters than our LongForm models.
OPT Zhang et al. (2022) is an open-source decoder-only transformer LM, trained on diverse English corpora. Due to hardware limitations that make long text inference from LLMs time-consuming, we use the 30B variant of OPT. T0++ Sanh et al. (2022) is an LM-adapted variant of T5-11B, instruction-tuned on PromptSource Bach et al. (2022), a dataset of $>$12 million examples from existing datasets. Tk-Instruct Wang et al. (2022b) is another instruction-tuned LM based on an LM-adapted version of T5-11B. The training data is a subsample of NIv2, which includes 757 tasks and over 75,000 examples. As we are performing zero-shot long text generation via instruction, we use the definition-only version of Tk-Instruct.
Flan-T5 Chung et al. (2022) is an instruction-tuned LM built on the LM-adapted version of T5-11B. The training data includes a variety of tasks and prompts from multiple sources, including Muffin Wei et al. (2022), T0-SF Sanh et al. (2022), and NIv2 Wang et al. (2022b), comprising more than 14 million examples in total. Alpaca is an instruction-tuned LM built on LLaMA-7B Touvron et al. (2023) by using a variation of Self-Instruct Wang et al. (2022a). As they did not release their finetuned LM, we finetune LLaMA-7B with their variation of Self-Instruct. [^1] https://github.com/tatsu-lab/stanford_alpaca/
@@ 4.3 Evaluation
We evaluate the performance of the baselines and our instruction-tuned LMs on different sizes and architectures. For generation, we perform nucleus sampling Holtzman et al. (2020) with $p=0.9$ for all LMs. We first evaluate them on the test set of LongForm-C; recall that it consists of a diverse set of NLP tasks and corpus examples. Additionally, we assess the generalization capabilities of our LMs on a set of long text generation datasets that were not seen during finetuning. These tasks include recipe generation Bień et al. (2020) from a given set of ingredients, story generation for a given prompt from Writing Prompts (WP) subreddit, and long form question answering, ELI5 Fan et al. (2019). We sample 250 examples from each task, resulting in 750 examples in total. Although these datasets are not part of LongForm-C, recipe generation is part of the instruction tuning data of two baselines, Tk-Instruct and Flan-T5. Next we comparatively analyze long-form generation and language understanding abilities. We compare the performance of LMs on out-of-domain long text generation tasks (i.e., RGen, ELI5, WP) and MMLU Hendrycks et al. (2021). We compare LMs finetuned with the LongForm-C dataset and the FLAN collection, [^2] Wang et al. (2023a)’s FLAN version Longpre et al. (2023) which has a strong performance on MMLU Chung et al. (2022). We also compare multilingual text generation capabilities. We reformulate the multilingual summarization corpus, MLSUM Scialom et al. (2020). Specifically, we provide an instruction in French, German, Spanish, and Russian to generate news articles in those languages for given titles. We utilize the translation of the following instruction: “Write a news article for the given title: X” where X is the title and the output is the full news article. We collect 100 samples for each language and then evaluate on news article generation. As current metrics in text generation have limited capabilities in evaluating long text generation Celikyilmaz et al. (2020), we choose METEOR Banerjee and Lavie (2005) as our main metric as it exhibits higher human correlation Sharma et al. (2017); Chen et al. (2022). However, we also report BLEU Papineni et al. (2002), ROUGE Lin (2004), and self-BLEU Zhu et al. (2018) scores in the Appendix (§A), which exhibit similar patterns.
@@ 5 Results
We now present a comprehensive evaluation of our LongForm models compared to the baselines on LongForm-C test, recipe generation, long form question answering, short story generation, NLU and multilingual news generation.
@@ 5.1 LongForm-C Results
We begin by comparing baseline LMs and LongForm models on the test set of LongForm-C. As presented in Table 2, all LongForm models demonstrate clear improvements in performance compared to the prior instruction-tuned and raw LMs. In particular, LongForm-T5-XL, with 3B parameters, outperforms all instruction-tuned LMs based on T5 (i.e., T0++, Tk-Instruct, Flan-T5), even with 11B parameters, across all subtasks. This suggests the effectiveness of LongForm-C in finetuning LMs to follow instructions for long text generation. Moreover, among our LMs, we see that LongForm-LLaMA-7B outperforms LongForm-OPT-6.7B, despite both having similar number of parameters, which aligns with findings from Touvron et al. (2023). Additionally, we observe that LongForm-T5-XL has comparable or better results than auto-regressive variants, particularly on text generation (i.e., BEA-GEC, NIv2, BigBench). However, a potential limitation of T5 variants is the absence of newline tokens in their vocabulary, which may impact the readability of the generated output, particularly for long text generation. Among baseline LMs, Alpaca, the only instruction tuned LM without reformulated NLP tasks, achieves the best overall performance. However, it still performs worse than our LongForm models, potentially due to the noise in its training data, Self-Instruct. Furthermore, OPT-30B, despite having the highest number of parameters, fails to follow instructions as effectively as instruction-tuned LongForm-T5-XL and LongForm-OPT-2.7B, which are ten times smaller (see Figure 4). This observation is consistent with prior findings that demonstrate the effectiveness of instruction tuning in smaller LMs Ouyang et al. (2022).
@@ 5.2 Out-of-domain generalization
Table 3: METEOR scores for out-of-domain datasets. LongForm models outperform prior instruction-tuned LMs on recipe generation (RGen), long-form question answering (ELI5) and short story generation (WritingPrompts/WP). † indicates that the training data of Tk-Instruct and FLAN-T5 contain RGen.
To evaluate the generalization capability of LongForm models, we evaluate on three diverse datasets: recipe generation from ingredients (RGen) Bień et al. (2020), long form question answering (ELI5) Fan et al. (2019), and short story generation from the WritingPrompts subreddit (WP). We note that the LongForm models are not trained on any of these three datasets while the training set of Tk-Instruct and Flan-T5 includes RGen.
Table 3 shows that the LongForm models consistently outperform baseline LMs in all tasks, with LongForm-LLaMA-7B achieving a 35% higher relative METEOR score than the best baseline LM, Alpaca-LLaMA-7B. While they both have the same underlying LM, our LM achieves higher scores than Alpaca. This can be attributed to the diverse and human-written nature of LongForm-C compared to Alpaca’s fully synthetic examples. We present qualitative examples from the best performing LongForm model and the best baseline LM, Alpaca, for each task in Figure 5 in the Appendix (§D). These examples illustrate that LongForm models demonstrate better factual accuracy, instruction-following, long form text planning capabilities than prior instruction-tuned LMs.
5.3 Language Understanding and Generation
We evaluate how well LongForm-C helps to improve natural language understanding (NLU) in addition to generation tasks. For this purpose, we compare LongForm-C with a strong baseline, the FLAN collection Longpre et al. (2023), which performs well on NLU. We finetune the same base LM, LLaMA-7B, and compare the two LMs and another LM finetuned with both datasets on OOD generation tasks (i.e., recipe generation, ELI5 question answering, and story generation) and 5-shot MMLU performance. Table 4 shows that LLaMA finetuned with LongForm-C outperforms the baseline on language generation (+10.6). However, we see that the baseline is slightly better on MMLU. Incorporating LongForm-C with FLAN achieves the best MMLU performance (+2.3) however it does not achieve as good a performance as the LM with only LongForm-C on NLG tasks (-3.2).
@@ 5.3 Language Understanding and Generation
Table 4: LongForm-C outperforms FLAN in NLG tasks, achieving the best METEOR score on a combination of RGen, ELI5, and WP. Although FLAN performs slightly better on MMLU, integrating it with LongForm-C improves 5-shot MMLU accuracy by 2.3 points over the baseline.
5.4 Multilingual News Generation
@@ 5.5 Data Evaluation
Our instruction generation (for reverse instructions) relies on LLM generation. To ensure its quality, we conduct a human evaluation. We randomly select 100 corpus examples from LongForm-C before adding length information and provide them to Mechanical Turk annotators. Annotators are asked to determine whether the generated instruction is relevant to the given document or not. 97 out of 100 instructions are found to be relevant. This high degree of relevance suggests that combining instruction generation with corpus examples can potentially improve the effectiveness of LLMs for instruction-following compared to the Self-Instruct dataset Wang et al. (2022a), which relies on LLMs to generate tasks, instructions, and outputs, and achieves a validity of 54%.
We conducted an additional manual analysis of a corpus subset of LongForm-C to identify areas for improvement in corpus mining and instruction tuning, providing valuable insights for future work. We present qualitative examples in Table 10 in the Appendix. Our analysis shows that the failure cases mostly stem from the detailed nature of the corpus texts such as news. While reverse instructions is capable of generating relevant instructions, the corpus examples can sometimes include extensive content. This tendency can lead LongForm models to generate long responses even for simple questions that require concise answers. This issue occurred infrequently, as supported by the overall performance of LongForm. However, this is still an important insight for future work in this area.
Avg. DE ES FR RU
T0++ 1.8 1.6 2.4 2.5 0.7
Tk-Instruct 0.8 0.8 1.1 1.0 0.5
Flan-T5 2.4 2.5 2.8 3.4 0.7
Alpaca-LLaMA-7B 4.9 7.9 7.0 4.1 0.7
OPT-30B 4.5 12.7 3.5 1.3 0.5
LongForm-T5-XL 6.9 5.2 9.0 9.1 4.2
LongForm-OPT-6.7B 9.5 7.3 12.2 15.3 3.1
LongForm-LLaMA-7B 9.7 8.9 10.5 14.8 4.7
Table 5: Evaluation of multilingual news generation via METEOR.
@@ 6 Conclusion
The proposed datasets and models mainly focus on long text generation and have limitations regarding structured prediction tasks in NLP. Additionally, we observe that LongForm models may present hallucination problems similar to those found in LLMs.
Appendix A Evaluation with Different Metrics
@@
Appendix B Multilingual Prediction

### v3

@@ 4.1 LongForm Models
We create three LongForm models by finetuning language models on the LongForm-C dataset to assess their effectiveness on instruction-following text generation: LongForm-T5-3B, based on T5-3B Raffel et al. (2020); Lester et al. (2021), an encoder-decoder LM, and two autoregressive LMs: LongForm-OPT-6.7B, based on OPT-6.7B Zhang et al. (2022), and LongForm-LLaMA-7B, based on LLaMA-7B Touvron et al. (2023a). We share the details of finetuning in the Appendix (§A).
@@ 4.2 Baselines
We compare our instruction-tuned LMs with other LMs that are instruction-tuned on other datasets. These LMs are based on either T5-11B Raffel et al. (2020); Lester et al. (2021) or LLaMA-7B Touvron et al. (2023a). We also compare our LMs with a raw large LM, OPT-30B Zhang et al. (2022). Notably, all LMs have the same or higher number of parameters than our LongForm models.
OPT Zhang et al. (2022) is an open-source decoder-only transformer LM, trained on diverse English corpora. Due to hardware limitations that make long text inference from LLMs time-consuming, we use the 30B variant of OPT.
T0++ Sanh et al. (2022) is an LM-adapted variant of T5-11B, instruction-tuned on PromptSource Bach et al. (2022), a dataset of $>$12 million examples from existing datasets.
Tk-Instruct Wang et al. (2022) is another instruction-tuned LM based on an LM-adapted version of T5-11B. The training data is a subsample of NIv2, which includes 757 tasks and over 75,000 examples. As we are performing zero-shot long text generation via instruction, we use the definition-only version of Tk-Instruct.
Flan-T5 Chung et al. (2024) is an instruction-tuned LM built on the LM-adapted version of T5-11B. The training data (total size 14M) includes tasks and prompts from multiple sources, including Muffin Wei et al. (2022), T0-SF Sanh et al. (2022), and NIv2 Wang et al. (2022).
Alpaca is an instruction-tuned LM built on LLaMA-7B Touvron et al. (2023a) by using a variation of Self-Instruct Wang et al. (2023b). As they did not release their finetuned LM, we finetune LLaMA-7B with their variation of Self-Instruct.111github.com/tatsu-lab/stanford_alpaca/
@@ 4.3 Evaluation
We evaluate the performance of the baselines and our instruction-tuned LMs on different sizes and architectures. For generation, we perform nucleus sampling Holtzman et al. (2020) with $p=0.9$ for all LMs. We first evaluate them on the test set of LongForm-C; recall that it consists of a diverse set of NLP tasks and corpus examples.
Additionally, we assess the generalization capabilities of our LMs on a set of long text generation datasets that were not seen during finetuning. These tasks include recipe generation Bień et al. (2020) from a given set of ingredients, story generation for a given prompt from Writing Prompts (WP) subreddit, and long form question answering, ELI5 Fan et al. (2019). We sample 250 examples from each task, resulting in 750 examples in total. Although these datasets are not part of LongForm-C, recipe generation is part of the instruction tuning data of two baselines, Tk-Instruct and Flan-T5.
Next we analyze long-form generation and language understanding abilities. We compare the performance of LMs on out-of-domain long text generation tasks (i.e., RGen, ELI5, WP) and MMLU Hendrycks et al. (2021). We compare LMs finetuned with the LongForm-C dataset and the FLAN collection, [^2] Wang et al. (2023a)’s FLAN version Longpre et al. (2023) which has a strong performance on MMLU Chung et al. (2024). We also extend this analysis on more recent and powerful LLMs. In addition to LLaMA-7B, we compare FLAN and LongForm-C datasets on LLaMA 2-7B (Touvron et al., 2023b) and Mistral 7B (Jiang et al., 2023).
We also compare on multilingual text generation. We reformulate the multilingual summarization corpus MLSUM Scialom et al. (2020) by providing an instruction in French, German, Spanish, Russian to generate news articles in those languages for given titles. We utilize the translation of the following instruction: “Write a news article for the given title: X” where X is the title and the output is the full news article. We collect 100 samples for each language and then evaluate on news article generation. We report results in §C in Appendix.
As current metrics in text generation have limited capabilities in evaluating long text generation Celikyilmaz et al. (2020), we choose METEOR Banerjee and Lavie (2005) as our main metric as it exhibits higher human correlation Sharma et al. (2017); Chen et al. (2022). However, we also report BLEU Papineni et al. (2002), ROUGE Lin (2004), and self-BLEU Zhu et al. (2018) scores in the Appendix (§B), which exhibit similar patterns.
@@ 5 Results
We now present a comprehensive evaluation of our LongForm models compared to the baselines on LongForm-C test, recipe generation, long form question answering, short story generation, NLU, and ablation of LongForm-C subsets.
@@ 5.1 LongForm-C Results
We begin by comparing baseline LMs and LongForm models on the test set of LongForm-C. As presented in Table 2, all LongForm models demonstrate clear improvements in performance compared to the prior instruction-tuned and raw LMs. In particular, LongForm-T5-XL, with 3B parameters, outperforms all instruction-tuned LMs based on T5 (i.e., T0++, Tk-Instruct, Flan-T5), even with 11B parameters, across all subtasks. This suggests the effectiveness of LongForm-C in finetuning LMs to follow instructions for long text generation.
Moreover, among our LMs, we see that LongForm-LLaMA-7B outperforms LongForm-OPT-6.7B, despite both having similar number of parameters, which aligns with findings from Touvron et al. (2023a). Additionally, we observe that LongForm-T5-XL has comparable or better results than auto-regressive variants, particularly on text generation (i.e., BEA-GEC, NIv2, BigBench). However, a potential limitation of T5 variants is the absence of newline tokens in their vocabulary, which may impact the readability of the generated output, particularly for long text generation.
Among baseline LMs, Alpaca, the only instruction tuned LM without reformulated NLP tasks, achieves the best overall performance. However, it still performs worse than our LongForm models, potentially due to the noise in its training data, Self-Instruct. Furthermore, OPT-30B, despite having the highest number of parameters, fails to follow instructions as effectively as instruction-tuned LongForm-T5-XL and LongForm-OPT-2.7B, which are ten times smaller (see Figure 4). This observation is consistent with prior findings that demonstrate the effectiveness of instruction tuning in smaller LMs Ouyang et al. (2022).
@@ 5.2 Out-of-domain Generalization
Table 3: METEOR scores for out-of-domain datasets. LongForm models outperform prior instruction-tuned LMs on recipe generation (RGen), long-form question answering (ELI5) and short story generation (WP). †: Tk-Instruct and FLAN-T5 are trained on RGen.
To evaluate the generalization capability of LongForm models, we evaluate on three diverse datasets: recipe generation (RGen) Bień et al. (2020), long form question answering (ELI5) Fan et al. (2019), and short story generation (WP). We note that the LongForm models are not trained on any of these three datasets while the training set of Tk-Instruct and Flan-T5 includes RGen.
Table 3 shows that the LongForm models consistently outperform baseline LMs in all tasks, with LongForm-LLaMA-7B achieving a 35% higher relative METEOR score than the best baseline LM, Alpaca-LLaMA-7B. While they both have the same underlying LM, our LM achieves higher scores than Alpaca. This can be attributed to the diverse and human-written nature of LongForm-C compared to Alpaca’s fully synthetic examples.
We present qualitative examples from the best performing LongForm model and the best baseline LM, Alpaca, in Figure 5 in the Appendix (§E). This illustrates that LongForm models have better factual accuracy, instruction-following, long form text planning than prior instruction-tuned LMs.
@@
LLaMA 2-7B FLAN 13.3 43.5
LongForm-C 19.3 44.7
FLAN+LongForm-C 17.8 46.2
Mistral 7B FLAN 18.9 57.7
LongForm-C 22.0 58.8
FLAN+LongForm-C 19.4 58.3
Table 4: LongForm-C consistently outperforms FLAN on NLG tasks (RGen, ELI5, WP) across all models. For LLaMA-7B, FLAN+LongForm-C achieves the best MMLU score, while LongForm-C outperforms FLAN in both NLG and NLU with LLaMA 2-7B and Mistral 7B: the improvements with reverse instructions are consistent and even more powerful with stronger models.
5.3 Language Understanding and Generation
We evaluate how much LongForm-C improves natural language understanding (NLU) in addition to NLG. To this end, we compare LongForm-C with the strong baseline FLAN Longpre et al. (2023), which performs well on NLU. We finetune base LMs using LongForm-C, FLAN and their combination, evaluating performance on OOD NLG (i.e., recipe generation, ELI5 question answering, and story generation) and 5-shot MMLU. We extend our evaluation to newer, more powerful models than LLaMA-7B: LLaMA 2-7B and Mistral 7B.
Table 4 shows that LLaMA-7B finetuned with LongForm-C outperforms the FLAN baseline in NLG (+10.6). Incorporating LongForm-C with FLAN achieves the best MMLU performance (+2.3), though the combination gets lower NLG scores than LongForm-C alone (-3.2).
Extending LLMs: Our analysis to more recent LLMs shows promising results. We observe that, with stronger models, LongForm-C shows better results than FLAN (both in NLU and NLG). With LLaMA 2-7B, LongForm-C again improves NLG (+6.0) while narrowing the gap in NLU, even surpassing FLAN’s MMLU score (+1.2). The FLAN+LongForm-C combination further improves MMLU (+2.7). With Mistral 7B, LongForm-C outperforms both FLAN and the combination, not only on NLG (+3.1, +2.6) but also on MMLU (+1.1, +0.5). These findings indicate that LongForm-C consistently improves NLG performance across LLMs, with its impact on NLU becoming more pronounced in stronger LLMs. This suggests that more natural datasets like LongForm-C may lead to better improvements in stronger LLMs than NLP reformulation datasets like FLAN.
5.4 Ablation of LongForm-C Subsets
We now conduct an ablation study to investigate the differential impact of the three distinct subsets of LongForm-C by finetuning LLaMA-7B: NLP Tasks (NLP), Structured Corpora (SC), and Reverse Instructions (RI).
Model Subset NLG NLU (MMLU)
LLaMA-7B NLP 13.6 35.4
NLP+SC 18.6 35.0
NLP+RI 19.9 35.5
NLP+RI+SC 19.7 35.2
Table 5: Subset ablation experiments to study the effect of different subsets in LongForm-C. Adding the Reverse Instructions subset further improves the performance both on NLG and NLU showing its capabilities while NLP and SC subsets show limited improvements.
We summarize the results in Table 5. Since the NLP tasks subset already includes examples with long form output, the model performs comparatively well on NLG, achieving better performance than the performance of FLAN (see Table 4), another NLP tasks dataset for instruction-tuning. We observe a trade-off between NLG (+5.0) and NLU (-0.4) performance when we include the SC subset in addition to NLP tasks. However, including the Reverse Instructions subset significantly improves NLG (+6.3) and also slightly improves NLU (+0.1). Despite the instructions in the SC subset being created directly from human annotation (StackExchange) or via minimal changes (WikiHow), the improvement by Reverse Instructions is substantially better and more consistent. We achieve the best results when the Reverse Instructions (RI) subset is included.
@@ 5.5 Data Evaluation
Our instruction generation (for reverse instructions) relies on LLM generation. To ensure its quality, we conduct a human evaluation. We randomly select 100 corpus examples from LongForm-C (without adding length information) and provide them to Mechanical Turk annotators. Annotators are asked to determine whether the generated instruction is relevant to the given document or not. 97 out of 100 instructions are found to be relevant. This high degree of relevance suggests that combining instruction generation with corpus examples can potentially improve the effectiveness of LLMs for instruction-following compared to the Self-Instruct dataset Wang et al. (2023b), which relies on LLMs to generate tasks, instructions, and outputs, and achieves a validity of 54%.
We conducted an additional manual analysis of a corpus subset of LongForm-C to identify areas for improvement in corpus mining and instruction tuning, providing valuable insights for future work. We present qualitative examples in Table 11 in the Appendix. Our analysis shows that the failure cases mostly stem from the detailed nature of corpus texts such as news. While reverse instructions is capable of generating relevant instructions, the corpus examples can sometimes include extensive content. This tendency can lead LongForm models to generate long responses even for simple questions that require concise answers. This issue occurred infrequently, as supported by the overall performance of LongForm. However, this is still an important insight for future work in this area.
@@ 6 Conclusion
The proposed datasets and models mainly focus on long text generation and may have limitations regarding structured prediction tasks in NLP with some LLMs. We observe that LongForm-C is a good complementary dataset and when combined with existing datasets, it further improves NLU performance. However, for more recent models, it outperforms previous datasets not only in NLG but also on structured prediction tasks. Additionally, we observe that LongForm models may present hallucination problems similar to those found in LLMs.
Acknowledgements
This work was funded by Deutsche Forschungsgemeinschaft (project SCHU 2246/14-1).
Appendix A Training Details
We include an end-of-instruction token, [EOI], between the instruction and the output of the autoregressive LMs, OPT and LLaMA variants, to differentiate between instruction and output. We employ two to four A6000 48GB GPUs for finetuning with a batch size of 32 to 64, and train with DeepSpeed, including bfloat16 training. The learning rate is determined by comparing the minimum validation loss during three epochs: 5e-5 for LongForm-T5-3B, 5e-6 for LongForm-OPT-6.7B and LongForm-LLaMA-7B. Apart from the three large LMs, we finetune four smaller OPT LMs (125M, 350M, 1.3B, 2.7B) with LongForm-C and release them to increase accessibility of instruction-tuned LMs. See §D for more details.
Appendix B Evaluation with Different Metrics
@@
Avg. DE ES FR RU
T0++ 1.8 1.6 2.4 2.5 0.7
Tk-Instruct 0.8 0.8 1.1 1.0 0.5
Flan-T5 2.4 2.5 2.8 3.4 0.7
Alpaca-LLaMA-7B 4.9 7.9 7.0 4.1 0.7
OPT-30B 4.5 12.7 3.5 1.3 0.5
LongForm-T5-XL 6.9 5.2 9.0 9.1 4.2
LongForm-OPT-6.7B 9.5 7.3 12.2 15.3 3.1
LongForm-LLaMA-7B 9.7 8.9 10.5 14.8 4.7
Table 9: Evaluation of multilingual news generation via METEOR.
Appendix C Multilingual News Generation